OpenStudy (anonymous):

Please walk me through this, I'm not sure how to do this one. :/ @jim_thompson5910

jimthompson5910 (jim_thompson5910):

We're using compound interest, but not continuously compounded interest since it only compounds 2 times per month, ie 2*12 = 24 times per year

jimthompson5910 (jim_thompson5910):

so we'll need this formula A = P*(1+r/n)^(n*t)

jimthompson5910 (jim_thompson5910):

A = final balance P = initial balance (aka principal) r = interest rate n = compounding frequency t = time in years

jimthompson5910 (jim_thompson5910):

so in this case, A = 2134.96 P = 1327.35 r = unknown (we're solving for this) n = 2*12 = 24 t = 12

jimthompson5910 (jim_thompson5910):

plug those values in and solve for r A = P*(1+r/n)^(n*t) 2134.96 = 1327.35*(1+r/24)^(24*12) r = ??

OpenStudy (anonymous):

r= 1.00 ? That's what I got. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

r = 1 implies you have a 100% interest rate

jimthompson5910 (jim_thompson5910):

which is just beyond crazy

jimthompson5910 (jim_thompson5910):

one sec and I'll write out how to solve for r

OpenStudy (anonymous):

I solved for "r" and that's what the calculator gave me. It was more like 1.00165 ...., but still. :/ ~Alright, thanks! That'd be great! :DD

jimthompson5910 (jim_thompson5910):

A = P*(1+r/n)^(n*t) 2134.96 = 1327.35*(1+r/24)^(24*12) 2134.96/1327.35 = (1+r/24)^(24*12) 1.60843786 = (1+r/24)^(288) (1.60843786)^(1/288) = 1+r/24 1.00165158 = 1+r/24 1.00165158-1 = r/24 0.00165158 = r/24 r/24 = 0.00165158 r = 24*0.00165158 r = 0.03963792 So the interest rate is around 3.96% which rounds to 4% Usually, most interest rates are less than 10%. The higher interest rates like 20% or 30% (or sometimes even 40%) are for very risky things like credit cards.

jimthompson5910 (jim_thompson5910):

0.3% is NOT the same as 3%

jimthompson5910 (jim_thompson5910):

I think that's where the mixup is

jimthompson5910 (jim_thompson5910):

it's 4.0% since that's the same as 4%

OpenStudy (anonymous):

Right; oh, I see the mixup. Thank you!!